# Searching for Exotrojans in Pulsar Binary Systems

### v1

@@ 4.4 Librating Trojan Prevalence
In general, for $N$ non-detections, $\eta_{\rm T,Lib}^{\rm UL}(N)=1-0.05^{1/(N+1)}$. We have, $\eta_{\rm T,Lib}^{\rm UL}(N)=|\ln{(0.05)}|/N+\mathcal{O}(1/N^{2})\approx 3/N$. Therefore, $\eta_{\rm T,Lib}^{\rm UL}(N)$ scales like $1/N$, so more non-detections constraint $\eta_{\rm T,Lib}$ further, but at a decreasing rate.
5 Discussion and Conclusions
@@
This study presents the first search for pulsar-bound exotrojans. Trojans in general are difficult to detect through methods that rely exclusively on the reflex motion of their host star because these effects are degenerate with those of the dominant co-orbital. Therefore, standard pulsar exoplanet searches are blind to co-orbital bodies. Trojans instead require explicit searches, due to their unique, subtle effects on the pulse times of arrival. A librating Trojan detection would improve constraints placed on the dominant companion mass and orbital inclination, giving good motivation for targeted Trojan searches.
@@
We used two distinct methods to place upper mass limits on Trojans in eight pulsar binary systems. One method combines radio pulsar timing with optical-band light curves in order to determine the difference in location between the fictitious combined companion and the actual dominant companion. Upon finding no significant difference, we concluded that the mass difference between the triangular Lagrange points can be no more than 8 MJ in the PSR J1641+8049 system. Our use of this method reaffirms the symmetry of black widow heating models. The other method requires no optical data but instead relies on the Trojan and dominant companion librating with respect to their stable configuration. Despite the highly precise pulse times of arrival in the NG15yr narrowband dataset, we find only weak evidence of librations in only two of the eight NG15yr systems studied here. We place $\sim 1\,\rm{M}_{\oplus}$ upper limits on the eight NG15yr systems assuming some properties of the pulsar population as well the hypothetical pulsar Trojan population (see Appendix B).
While we find signals consistent with librations in the PSR J0023+0923 and PSR J1703-1903 systems, the signals are likely associated with nonphysical Trojan properties whose uncertainties spread into physically acceptable ones. However, if the signals are rather from Trojans, we find $m_{\rm T}\sim 1\,\rm{M}_{\rm J}$. Other determined properties, like $m_{\rm c}$ and $I$, of the PSR J0023+0923 system are at odds with D. Mata Sánchez et al. (2023), weakening any conclusion that the detected signal is Trojan-caused. Furthermore, $\mathcal{B}_{\xi}<1$ for both pulsar systems, meaning the detections of the Equation 3.2 timing model are almost certainly false positives with respect to Trojan detection. We suspect that the true origin is due to complex and chaotic mass transfers in black widow systems, due to the fact that PSR J0023+0923 has known orbital variability (A.-S. Bak Nielsen et al., 2020; G. Agazie et al., 2023) and PSR J1703-1903 experiences radio eclipses for roughly 30% of its orbit (G. Agazie et al., 2023). Indeed, the timing of both of these pulsars requires five binary frequency time derivatives ($\rm{d}n_{\rm b}/\rm{d}t,\,\rm{d}^{2}n_{\rm b}/\rm{d}t^{2},\,...,\,\rm{d}^{5}n_{\rm b}/\rm{d}t^{5}$; G. Agazie et al. 2023). We also identify statistically significant red noise in PSR J1703-1903. Additionally, G. Agazie et al. (2023) find that the parallax determination of PSR J0023+0923 is covariant with the binary frequency derivatives, further complicating its stable timing. That being said, as to why a model with two frequencies straddling the orbital frequency by the same margin ($\nu$; see Equation A19) is preferred over a simpler type of “noise” is unclear.
@@
In any case, the timing residuals of PSR J0023+0923 possibly favor ($\mathcal{B_{T}}=11.5(2)$) Equation 3.2 being included with the original timing model, and the residuals significantly prefer ($\mathcal{B_{T}}=5700(300)$) of the same for PSR J1705-1903. Therefore, the status of these two pulsars in future pulsar timing array datasets may be called into question. Notably, the MeerKAT Pulsar Timing Array has already removed PSR J1705-1903 from its array in its most recent data release (M. T. Miles et al., 2025). While C. Bochenek et al. (2015) show that including up to four orbital frequency derivatives does not affect a pulsar’s sensitivity to gravitational waves, the Trojan libration model has not been applied to any pulsar before this work. If these pulsars are to be included in future pulsar timing array datasets, such as the forthcoming NANOGrav 20-year dataset, more noise modeling or analysis may be required, which may weaken the effectiveness of these pulsars in detecting the gravitational wave background. Conversely, including this new source of “noise” may improve their effectiveness as compared to where the pulsars currently stand.
@@
There are several clear paths to continuing the pulsar Trojan search. There exist myriad light curves of pulsar binary systems, all of which can have their times of superior conjunction compared to those predicted by pulsar timing. Furthermore, if several light curves are obtained on the same pulsar with the explicit goal of lowering the uncertainty in the time of superior conjunction, libration-independent constraints on the pulsar Trojan population will lower in tandem. In terms of the libration method, plenty of TOA data sets are available for analysis, with several high-quality ones in the making. Currently available high-cadence datasets include those from the the European Pulsar Timing Array (J. Antoniadis et al., 2023), the MeerKAT Pulsar Timing Array (M. T. Miles et al., 2025), and the Parkes Pulsar Timing Array (A. Zic et al., 2023). Gamma-ray pulsar timing arrays (e.g., F.-L. Collaboration*†, 2022; M. Kerr et al., 2024) present yet another way to search for Trojan librations. Because this work was done on the NG15yr dataset, the NANOGrav 20-year dataset will allow for the quickest updates on the Trojan mass constraints or may even reveal a detection.
@@
Our analysis reveals that pulsar Trojans detectable through their librations are likely somewhat rare, with there being no more than one for every three pulsars that satisfy Equation 1. One explanation is that pulsar Trojans are strongly damped into their stable configuration point, lowering the libration amplitude. This seems plausible since pulsar Trojans would have to form from a dense debris or accretion disk. Alternatively, noting that $\mathcal{T}\propto z\cdot m_{\rm T}$, the pulsar Trojan masses may simply be too low to detect for their given distribution in libration amplitude. Another explanation is that the conditions amenable to Trojan formation are also poor environs for those same Trojans. Mass transfer and disk dynamics in pulsar binary systems are clearly turbulent, and so any interloping Trojan may find itself ejected from the system or torn apart. Lastly, conditions in pulsar binaries may never be friendly to Trojan formation to begin with. Future pulsar Trojan searches, as discussing above, will aid in lowering the upper limit on the librating pulsar Trojan prevalence, and will also constrain the non-librating pulsar Trojan occurrence rate through the superior conjunction method.
@@
Our work here not only represents the first search for Trojans in pulsar binary systems, but also, along with L. Dey et al. (2025), showcases a rigorous and thorough way to analyze deterministic signals in pulse TOAs. The search procedure can be extended to ordinary pulsar exoplanets, pulsar exomoons (or, equivalently, in terms of TOA residuals, planets in the S-type orbital configuration), and even massive bodies in our own Solar System. Such a Solar System body would be correlated across all pulsars, because the location of the Solar System barycenter would shift unexpectedly in time. This is the so-called ”Earth term” and can be used to, for example, detect unknown massive Trojans (of any of the planets), discover unknown primordial black holes that orbit the sun, or constrain the mass of the known objects in the solar system.
@@
The NG15yr narrowband data set is available at https://zenodo.org/records/14773896
@@
J.D.T. led the project, extended the libration theory of A. Leleu et al. (2015) as relevant for black widow systems, produced the figures, conducted the analysis, and prepared the text with input from E.F., L.D., S.Z., A.Y.K., and J.G. L.D. provided the original code infrastructure for the use of ENTERPRISE and PTMCMC before the FFO-to-Trojan transition and offered key insights into the Bayesian inference used. S.Z. and A.Y.K. carried out the optical-band light curve analysis in determining the optical times of superior conjunction and wrote the corresponding Section 2.1. A.Y.K. provided the optical light curve data and contributed useful discussions regarding PSR J1641+8049. J.G. provided essential high-performance computing support. J.G., P.B.D., M.K., D.J.N., and D.L.K. contributed insightful discussions. E.F., J.G., P.B.D., M.K., D.J.N., D.L.K., and the remaining authors contributed to the collection and analysis of the NANOGrav 15 yr dataset.
@@
We thank Reynier Squillace for helpful comments. Computational resources were provided by the Link HPC cluster and cyber-infrastructure, which is maintained by the Center for Gravitational Waves and Cosmology at West Virginia University and funded in part by NSF IIA-1458952 & NSF PHY-2020265. Additional computational resources were provided by the WVU Research Computing Thorny Flat HPC cluster, partly funded by NSF OAC-1726534. This research was made possible by the NASA West Virginia Space Grant Consortium, Grant # 80NSSC25M7079, which provided individual support to J.D.T. This work has been carried out as part of the NANOGrav collaboration, which receives support from the National Science Foundation (NSF) Physics Frontiers Center award numbers 1430284 and 2020265. The Arecibo Observatory is a facility of the NSF operated under a cooperative agreement (No. AST-1744119) by the University of Central Florida (UCF) in alliance with Universidad Ana G. Méndez (UAGM) and Yang Enterprises (YEI), Inc. The Green Bank Observatory is a facility of the NSF operated under a cooperative agreement by Associated Universities, Inc. The National Radio Astronomy Observatory is a facility of the NSF operated under a cooperative agreement by Associated Universities, Inc. P.R.B. is supported by the Science and Technology Facilities Council, grant number ST/W000946/1. P.R.B. is supported by the Science and Technology Facilities Council, grant number ST/W000946/1. H.T.C. acknowledges funding from the U.S. Naval Research Laboratory. Pulsar research at UBC is supported by an NSERC Discovery Grant and by CIFAR. K.C. is supported by a UBC Four Year Fellowship (6456). M.E.D. acknowledges support from the Naval Research Laboratory by NASA under contract S-15633Y. T.D. and M.T.L. received support by an NSF Astronomy and Astrophysics Grant (AAG) award number 2009468 during this work. E.C.F. is supported by NASA under award number 80GSFC24M0006. D.C.G. is supported by NSF Astronomy and Astrophysics Grant (AAG) award #2406919. D.R.L. and M.A.M. are supported by NSF #1458952. M.A.M. is supported by NSF #2009425. The Dunlap Institute is funded by an endowment established by the David Dunlap family and the University of Toronto. T.T.P. acknowledges support from the Extragalactic Astrophysics Research Group at Eötvös Loránd University, funded by the Eötvös Loránd Research Network (ELKH), which was used during the development of this research. H.A.R. is supported by NSF Partnerships for Research and Education in Physics (PREP) award No. 2216793. S.M.R. and I.H.S. are CIFAR Fellows. Portions of this work performed at NRL were supported by ONR 6.1 basic research funding. S.Z. and A.K. thank Daniel Mata Sanchez for providing the optical data. SZ acknowledges the DGAPA-PAPIIT grant IN119323. AK acknowledges the DGAPA-PAPIIT grant IA105024.
@@ Appendix A Effect of Librations on Pulsar Times of Arrival
Here, we largely follow the calculations of A. Leleu et al. (2015) except we include the $\mathcal{O}(\nu/n_{\rm{b}})$ term that cannot be ignored for the pulsar companion masses of interest in our search. Many of the variables are defined in Section 3.2. Referencing Figure 2, we have the Trojan and companion librating with respect to each other, with the libration angle given by Equation 6. The mass-averaged longitude, $\lambda=n_{\rm b}\tilde{t}+\lambda_{0}$, follows the mean motion and is unaffected by the librations. The angles of the Trojan and companion with respect to $\lambda$ are $\delta\zeta$ and $(1-\delta)\zeta$, respectively. Therefore, as P. Robutel et al. (2011) showed,
@@
The semi-major axes of the Trojan and the companion in the pulsar-centric reference frame also oscillate with respect to the mass-averaged semi-major axis, $a$, and are given by P. Robutel et al. (2011) as
@@
is the term that is completely degenerate with that of a single companion and {widetext}
$\boldsymbol{r}_{\rm L}=\mu az\delta(1-\delta)\left(i\cos{(\nu\tilde{t}+\varphi)}+\frac{2\nu}{3n_{\rm b}}\sin{(\nu\tilde{t}+\varphi)}\right)e^{i\pi/3}e^{i(\lambda+\delta\pi/3)}+\mathcal{O}\left(\frac{\nu}{n_{\rm b}}z^{2}\right)$ (A10)
represents non-degenerate terms due to librations. By convention, we place the Earth such that it lies in the $xz$-plane (see Figure 2) and makes an angle $I$ with the $z$-axis. Therefore, by Equation 4,
@@
$\Re(\boldsymbol{r}_{\rm f})\frac{\sin{I}}{c}=$ $\frac{\mu a\sin{I}}{c}\sqrt{1-\delta(1-\delta)}$ (A12)
$\times\cos{\left[n\tilde{t}+\lambda_{0}+\delta\frac{\pi}{3}-\arctan{\left(\frac{\delta\sqrt{3}}{2-\delta}\right)}\right]}.$
This is already included in the timing model (Section 3.2.1), and so we can exploit the fact that $T_{\rm{asc}}$ and $K=\alpha\sqrt{1-\delta(1-\delta)}=\mu a\sin{I}\sqrt{1-\delta(1-\delta)}/c$ are already known. Without prior knowledge of a Trojan,
@@
To zeroth order in the orbital eccentricity, C. Lange et al. (2001) defined $T_{\rm{asc}}$ such that $\Delta\tau_{2}=K\sin{(\Phi)}=K\sin{(n_{\rm b}\tilde{t})}$, where $\Delta\tau_{2}$ is the time delay from an ordinary two-body system. $\phi=\Phi-\pi/2$ is shown in Figure 2. The $\pi/2$ phase difference arises from how C. Lange et al. (2001) reference $\Phi$ to the negative $y$-axis of our convention. Matching the phases of $\Delta\tau_{2}$ and $\Re(\boldsymbol{r}_{\rm f})\sin{(I)/c}$ gives
@@
Note that $\phi$ and $\lambda$ are close but not identical, with
@@
Continuing Equation A11 with the non-degenerate term, and ignoring $\mathcal{O}(\nu z^{2}/n_{\rm b})$ effects, we have {widetext}
$\Re(\boldsymbol{r}_{\rm T})\frac{\sin{I}}{c}=\frac{\mu a\sin{I}}{c}z\delta(1-\delta)\left[\cos{(\nu\tilde{t}+\varphi)}\cos{\left(n\tilde{t}+\lambda_{0}+\delta\frac{\pi}{3}+\frac{5\pi}{6}\right)}+\frac{2\nu}{3n_{\rm b}}\sin{(\nu\tilde{t}+\varphi)}\cos{\left(n\tilde{t}+\lambda_{0}+\delta\frac{\pi}{3}+\frac{\pi}{3}\right)}\right].$ (A17)
@@
is defined for simplification. Combining Equations A15 and A18 gives Equation 9. The pulsar timing residuals that are not degenerate with the timing model are given by Equation A17 and simplified in Equation 3.2.
We can write Equation 3.2 in a way that shows the periodicity involved directly, and represents how it would appear in the frequency domain. Namely,
@@
where $\sum_{\pm}$ denotes the sum over the plus and minus terms (e.g., $\sum_{\pm}(a\pm b)=(a+b)+(a-b)$). Thus, in frequency space, we expect to see a double peak around $n_{\rm b}$ at $n_{\rm b}\pm\nu$, where the $n_{\rm b}-\nu$ peak is stronger. A. Leleu et al. (2015) identify this double peak structure in their Figure 12. We conducted three-body simulations in the integrator REBOUND (H. Rein & S. F. Liu, 2012; H. Rein & D. S. Spiegel, 2015) and verify that the $n_{\rm b}-\nu$ peak is stronger when $\mathcal{O}(\nu/n_{\rm b})$ term cannot be ignored due to $\mu$ being sufficiently high (see Equation 10).
We identify two corrections to the model in A. Leleu et al. (2015), even after we remove the $\mathcal{O}(\nu/n_{\rm b})$ term. The first correction stems from their Equation 30, line 2. Here, they list $i(1-\delta)$, which instead should be $i(\delta-1)$. This is because in their Equations 9 and 10, there is a minus sign in front of the $i(1-\delta)$ exponent. This error leads to an additional factor of $\sqrt{3}$ in their derivation of $S_{1}^{*}=S_{-1}^{*}$ (see their Equation 32), which are equivalent to our $\mathcal{T}/2$. We place a superscript asterisk (*) on variables in their notation to avoid confusion.

### v2

@@ IV.4 Librating Trojan Prevalence
In general, for $N$ non-detections, $\eta_{\rm T,Lib}^{\rm UL}(N)=1-0.05^{1/(N+1)}$. We have $\eta_{\rm T,Lib}^{\rm UL}(N)=|\ln{(0.05)}|/N+\mathcal{O}(1/N^{2})\approx 3/N$. Therefore, $\eta_{\rm T,Lib}^{\rm UL}(N)$ scales like $1/N$, so more non-detections constraint $\eta_{\rm T,Lib}$ further, but at a decreasing rate.
V Discussion and Conclusions
@@
This study presents the first search for pulsar-bound exotrojans. Trojans in general are difficult to detect through methods that rely exclusively on the reflex motion of their host star because these effects are degenerate with those of the dominant co-orbital. Therefore, standard pulsar exoplanet searches cannot probe for co-orbital bodies. Trojans instead require explicit searches, due to their unique, subtle effects on the pulse times of arrival. A librating Trojan detection would improve constraints placed on the secondary mass and orbital inclination, giving good motivation for targeted Trojan searches.
@@
We used two distinct methods to place upper mass limits on Trojans in eight pulsar binary systems. One method combines radio pulsar timing with optical-band light curves in order to determine the difference in location between the fictitious combined companion and the actual companion. Upon finding no significant difference, we concluded that the mass difference between the triangular Lagrange points can be no more than 8 MJ in the PSR J1641+8049 system. Our use of this method reaffirms the symmetry of black widow heating models. The other method requires no optical data but instead relies on the Trojan and secondary librating with respect to their stable configuration. Despite the highly precise pulse times of arrival in the NG15yr narrowband dataset, we find only weak evidence of librations in only two of the eight NG15yr systems studied here. We place $\sim 1\,\rm{M}_{\oplus}$ upper limits on the eight NG15yr systems assuming some properties of the pulsar population as well the hypothetical pulsar Trojan population (see Appendix B).
[…]
@@
In any case, the timing residuals of PSR J0023+0923 possibly favor ($\mathcal{B_{T}}=11.5(2)$) Equation II.2 being included with the original timing model, and the residuals significantly prefer ($\mathcal{B_{T}}=5700(300)$) of the same for PSR J1705-1903. Therefore, the status of these two pulsars in future pulsar timing array datasets may be called into question. Notably, the MeerKAT Pulsar Timing Array has already removed PSR J1705-1903 from its array in its most recent data release (Miles et al., 2025). While Bochenek et al. (2015) show that including up to four orbital frequency derivatives does not affect a pulsar’s sensitivity to gravitational waves, the Trojan libration model has not been applied to any pulsar before this work. If these pulsars are to be included in future pulsar timing array datasets, such as the forthcoming NANOGrav 20-year dataset, more noise modeling or analysis may be required, which may weaken the effectiveness of these pulsars in detecting the gravitational wave background. Conversely, including this new source of “noise” may improve their effectiveness as compared to where the pulsars currently stand.
@@
There are several clear paths to continuing the pulsar Trojan search. There exist myriad light curves of pulsar binary systems, all of which can have their times of superior conjunction compared to those predicted by pulsar timing. Furthermore, if several light curves are obtained on the same pulsar with the explicit goal of lowering the uncertainty in the time of superior conjunction, libration-independent constraints on the pulsar Trojan population will lower in tandem. In terms of the libration method, plenty of TOA data sets are available for analysis, with several high-quality ones in the making. Currently available high-cadence datasets include those from the the European Pulsar Timing Array (Antoniadis et al., 2023), the MeerKAT Pulsar Timing Array (Miles et al., 2025), and the Parkes Pulsar Timing Array (Zic et al., 2023). Gamma-ray pulsar timing arrays (e.g., Collaboration*†, 2022; Kerr et al., 2024) present yet another way to search for Trojan librations. Because this work was done on the NG15yr dataset, the NANOGrav 20-year dataset will allow for the quickest updates on the Trojan mass constraints or may even reveal a detection.
@@
Our analysis reveals that pulsar Trojans detectable through their librations are likely somewhat rare, with there being no more than one for every three pulsars that satisfy Equation 1. One explanation is that pulsar Trojans are strongly damped into their stable configuration point, lowering the libration amplitude. This seems plausible since pulsar Trojans would have to form from a dense debris or accretion disk. Alternatively, noting that $\mathcal{T}\propto z\cdot m_{\rm T}$, the pulsar Trojan masses may simply be too low to detect for their given distribution in libration amplitude. Another explanation is that the conditions amenable to Trojan formation are also poor environs for those same Trojans. Mass transfer and disk dynamics in pulsar binary systems are clearly turbulent, and so any interloping Trojan may find itself ejected from the system or torn apart. Lastly, conditions in pulsar binaries may never be friendly to Trojan formation to begin with. Future pulsar Trojan searches, as discussed above, will aid in lowering the upper limit on the librating pulsar Trojan prevalence, and will also constrain the non-librating pulsar Trojan occurrence rate through the superior conjunction method.
@@
Our work here not only represents the first search for Trojans in pulsar binary systems, but also, along with Dey et al. (2026), showcases a rigorous and thorough way to analyze deterministic signals in pulse TOAs. The search procedure can be extended to ordinary pulsar exoplanets, pulsar exomoons (or, equivalently, in terms of TOA residuals, planets in the S-type orbital configuration), and even massive bodies in our own Solar System. Such a Solar System body would be correlated across all pulsars, because the location of the Solar System barycenter would shift unexpectedly in time. This is the so-called “Earth term” and can be used to, for example, detect unknown massive Trojans of any of the planets, discover unknown primordial black holes that orbit the sun, or constrain the mass of the known objects in the Solar System.
@@
The NG15yr narrowband data set is available at https://zenodo.org/records/14773896 (The NANOGrav Collaboration, 2025).
@@
J.D.T. led the project, extended the libration theory of Leleu et al. (2015) as relevant for black widow systems, produced the figures, conducted the analysis, and prepared the text with input from E.F., L.D., S.Z., A.Y.K., and J.G. L.D. provided the original code infrastructure for the use of ENTERPRISE and PTMCMC before the FFO-to-Trojan transition and offered key insights into the Bayesian inference used. S.Z. and A.Y.K. carried out the optical-band light curve analysis in determining the optical times of superior conjunction and wrote the corresponding Section III.1. A.Y.K. provided the optical light curve data and contributed useful discussions regarding PSR J1641+8049. J.G. provided essential high-performance computing support. J.G., P.B.D., M.K., D.J.N., D.L.K., and T.J.W.L. contributed insightful discussions. T.J.W.L. and D.R.L. gave helpful comments. E.F., J.G., P.B.D., M.K., D.J.N., D.L.K., D.R.L., and the remaining authors contributed to the collection and analysis of the NANOGrav 15 yr dataset.
@@
We thank Reynier Squillace and the anonymous reviewer for helpful comments. Computational resources were provided by the Link HPC cluster and cyber-infrastructure, which is maintained by the Center for Gravitational Waves and Cosmology at West Virginia University and funded in part by NSF IIA-1458952 & NSF PHY-2020265. Additional computational resources were provided by the WVU Research Computing Thorny Flat HPC cluster, partly funded by NSF OAC-1726534. This research was made possible by the NASA West Virginia Space Grant Consortium, Grant # 80NSSC25M7079, which provided individual support to J.D.T. This work has been carried out as part of the NANOGrav collaboration, which receives support from the National Science Foundation (NSF) Physics Frontiers Center award numbers 1430284 and 2020265. The Arecibo Observatory is a facility of the NSF operated under a cooperative agreement (No. AST-1744119) by the University of Central Florida (UCF) in alliance with Universidad Ana G. Méndez (UAGM) and Yang Enterprises (YEI), Inc. The Green Bank Observatory is a facility of the NSF operated under a cooperative agreement by Associated Universities, Inc. The National Radio Astronomy Observatory is a facility of the NSF operated under a cooperative agreement by Associated Universities, Inc. P.R.B. is supported by the Science and Technology Facilities Council, grant number ST/W000946/1. P.R.B. is supported by the Science and Technology Facilities Council, grant number ST/W000946/1. H.T.C. acknowledges funding from the U.S. Naval Research Laboratory. Pulsar research at UBC is supported by an NSERC Discovery Grant and by CIFAR. K.C. is supported by a UBC Four Year Fellowship (6456). M.E.D. acknowledges support from the Naval Research Laboratory by NASA under contract S-15633Y. T.D. and M.T.L. received support by an NSF Astronomy and Astrophysics Grant (AAG) award number 2009468 during this work. E.C.F. is supported by NASA under award number 80GSFC24M0006. D.C.G. is supported by NSF Astronomy and Astrophysics Grant (AAG) award #2406919. D.R.L. and M.A.M. are supported by NSF #1458952. M.A.M. is supported by NSF #2009425. The Dunlap Institute is funded by an endowment established by the David Dunlap family and the University of Toronto. T.T.P. acknowledges support from the Extragalactic Astrophysics Research Group at Eötvös Loránd University, funded by the Eötvös Loránd Research Network (ELKH), which was used during the development of this research. H.A.R. is supported by NSF Partnerships for Research and Education in Physics (PREP) award No. 2216793. S.M.R. and I.H.S. are CIFAR Fellows. Portions of this work performed at NRL were supported by ONR 6.1 basic research funding. S.Z. and A.K. thank Daniel Mata Sanchez for providing the optical data. SZ acknowledges the DGAPA-PAPIIT grant IN119323. AK acknowledges the DGAPA-PAPIIT grant IA105024.
@@ Appendix A Effect of Librations on Pulsar Times of Arrival
Here, we largely follow the calculations of Leleu et al. (2015) except we include the $\mathcal{O}(\nu/n_{\rm{b}})$ term that cannot be ignored for the pulsar companion masses of interest in our search. Many of the variables are defined in Section II.2, but we do repeat some details here for clarity. Referencing Figure 2, we have the Trojan and companion librating with respect to each other, with the libration angle given by $\zeta=z\cos{(\nu\tilde{t}+\varphi)}+\pi/3$ (Equation 12). The mass-averaged longitude, $\lambda=n_{\rm b}\tilde{t}+\lambda_{0}$, follows the mean motion and is unaffected by the librations. The angles of the Trojan and companion with respect to $\lambda$ are $\delta\zeta$ and $(1-\delta)\zeta$, respectively. Therefore, as Robutel et al. (2011) showed,
@@
The semi-major axes of the Trojan and the companion in the pulsar-centric reference frame also oscillate with respect to the mass-averaged semi-major axis, $a$, and are given by Robutel et al. (2011) as
@@
is the term that is completely degenerate with that of a single companion and
$\boldsymbol{r}_{\rm L}=$ $\mu az\delta(1-\delta)\left(i\cos{(\nu\tilde{t}+\varphi)}+\frac{2\nu}{3n_{\rm b}}\sin{(\nu\tilde{t}+\varphi)}\right)$
$\times e^{i\pi/3}e^{i(\lambda+\delta\pi/3)}+\mathcal{O}\left(\frac{\nu}{n_{\rm b}}z^{2}\right)$ (A10)
represents non-degenerate terms due to librations. By convention, we place the Earth such that it lies in the $\hat{x}\hat{z}-$plane (see Figure 2) and makes an angle $I$ with the $\hat{z}-$axis. Therefore, by Equation 4,
@@
$\Re(\boldsymbol{r}_{\rm f})\frac{\sin{I}}{c}=$ $\frac{\mu a\sin{I}}{c}\sqrt{1-\delta(1-\delta)}$
$\times\cos{\left[n\tilde{t}+\lambda_{0}+\delta\frac{\pi}{3}-\arctan{\left(\frac{\delta\sqrt{3}}{2-\delta}\right)}\right]}.$ (A12)
This is already included in the timing model (Section II.2.1), and so we can exploit the fact that $T_{\rm{asc}}$ and $K=\alpha\sqrt{1-\delta(1-\delta)}=\mu a\sin{I}\sqrt{1-\delta(1-\delta)}/c$ are already known. Without prior knowledge of a Trojan,
@@
To zeroth order in the orbital eccentricity, Lange et al. (2001) defined $T_{\rm{asc}}$ such that $\Delta\tau_{2}=K\sin{(\Phi)}=K\sin{(n_{\rm b}\tilde{t})}$, where $\Delta\tau_{2}$ is the time delay from an ordinary two-body system and $\Phi=n_{\rm b}\tilde{t}$. $\phi=\Phi-\pi/2$ is shown in Figure 2. The $\pi/2$ phase difference arises from how Lange et al. (2001) reference $\Phi$ to the negative $\hat{y}-$axis of our convention. Therefore,
$\phi_{0}=-\frac{\pi}{2}.$ (A15)
Matching the phases of $\Delta\tau_{2}$ and $\Re(\boldsymbol{r}_{\rm f})\sin{(I)/c}$ gives
@@
Therefore, $\phi$ and $\lambda$ are close but not identical, with
@@
Note that $\phi\equiv\lambda$ only when $\delta=0$, 1/2, or 1.
Continuing Equation A11 with the non-degenerate term, and ignoring $\mathcal{O}(\nu z^{2}/n_{\rm b})$ effects, we have
$\Re(\boldsymbol{r}_{\rm T})\frac{\sin{I}}{c}=$ $\frac{\mu a\sin{I}}{c}z\delta(1-\delta)$
$\times\Big[\cos{(\nu\tilde{t}+\varphi)}\cos{\left(n\tilde{t}+\lambda_{0}+\delta\frac{\pi}{3}+\frac{5\pi}{6}\right)}$
$+\frac{2\nu}{3n_{\rm b}}\sin{(\nu\tilde{t}+\varphi)}\cos{\left(n\tilde{t}+\lambda_{0}+\delta\frac{\pi}{3}+\frac{\pi}{3}\right)}\Big].$ (A18)
@@
is defined for simplification. Combining Equations A16 and A19 gives Equation 17. The pulsar timing residuals that are not degenerate with the timing model are given by Equation A and simplified in Equation II.2.
We can write Equation II.2 in a way that shows the periodicity involved directly, and represents how it would appear in the frequency domain. Namely,
@@
where $\sum_{\pm}$ denotes the sum over the plus and minus terms (e.g., $\sum_{\pm}(a\pm b)=(a+b)+(a-b)$). Thus, in frequency space, we expect to see a double peak around $n_{\rm b}$ at $n_{\rm b}\pm\nu$, where the $n_{\rm b}-\nu$ peak is stronger. Leleu et al. (2015) identify this double peak structure in their Figure 12. We conducted three-body simulations in the integrator REBOUND (Rein and Liu, 2012; Rein and Spiegel, 2015) and verify that the $n_{\rm b}-\nu$ peak is stronger when $\mathcal{O}(\nu/n_{\rm b})$ term cannot be ignored due to $\mu$ being sufficiently high (see Equation 13).
We identify two corrections to the model in Leleu et al. (2015), even after we remove the $\mathcal{O}(\nu/n_{\rm b})$ term. The first correction stems from their Equation 30, line 2. Here, they list $i(1-\delta)$, which instead should be $i(\delta-1)$. This is because in their Equations 9 and 10, there is a minus sign in front of the $i(1-\delta)$ exponent. This error leads to an additional factor of $\sqrt{3}$ in their derivation of $S_{1}^{*}=S_{-1}^{*}$ (see their Equation 32), which are equivalent to our $\mathcal{T}/2$. We place a superscript asterisk (*) on variables in their notation to avoid confusion.